I need to find out how many feet (or acres) are in this triangle
Side A = 1062 feet
Side B = 1026 feet
Base = 390 feet
Thank you

Answer by checkley71(8403)

SEEING AS THIS IS NOT A RIGHT TRIANGLE THE AREA FORMULA IS:
AREA=SQRT[S(S-A)(S-B)(S-C)] WHERE S=(A+B+C)/2
S=(1062+1026+390)/2=2478/2=1239
AREA=SQRT[1239(1239-1062)(1239-1026)(1239-390)]
AREA=SQRT[1239*177*213*845]
AREA=198,673 ANSWER.